# Regularity of vector fields with piecewise regular curl and divergence

### v2

@@ Abstract
We prove piecewise Sobolev regularity of vector fields that have piecewise regular curl and divergence, but may fail to be globally continuous. The main idea behind our approach is to employ recently developed parametrices for the curl-operator and the regularity theory of Poisson transmission problems. We conclude our work by applying our findings to the heterogeneous time-harmonic Maxwell equations with either a) impedance, b) natural or c) essential boundary conditions and providing wavenumber-explicit piecewise regularity estimates for these equations.
@@ 1 Introduction
For $\ell\in\mathbb{N}_{0}$ we consider a bounded $\mathcal{C}^{\ell+2}$-domain $\Omega$ with boundary $\Gamma:=\partial\Omega$ and a vector field ${\bf v}$ which has piecewise ${\bf H}^{\ell}$-regular curl and $\operatorname{H}^{\ell}$-regular divergence, but may jump across certain $\mathcal{C}^{\ell+2}$-regular and mutually disjoint surfaces $\mathcal{I}_{1},\ldots,\mathcal{I}_{r}$ in the interior of $\Omega$. We show that if ${\bf v}$ satisfies certain transmission conditions across these surfaces of discontinuity and boundary conditions on $\Gamma$, then ${\bf v}$ is already piecewise ${\bf H}^{\ell+1}$-regular, and the ${\bf H}^{\ell+1}$-norm of ${\bf v}$ can be controlled by its piecewise curl, its piecewise divergence and the boundary data.
@@
Let us mention that this work is the first paper of an upcoming trilogy of research articles. The second paper [18] in this trilogy builds on the subsequent Theorem 2.5 and Theorem 2.6 and proves wavenumber-explicit piecewise analyticity of solutions ${\bf u}$ of (1.1), provided that the geometry and all given coefficients and data are (piecewise) analytic. Finally, the third work [19] of this upcoming trilogy provides a wavenumber-explicit analysis of the $hp$-finite element method applied to Maxwell’s equations (1.1) in the presence of piecewise analytic coefficients. In particular, if $k$ is the wavenumber, $h$ is the local mesh-width and $p$ is the local polynomial degree in the $hp$-FEM approximation, the third paper will show that the $hp$-FEM is quasi-optimal under the scale-resolution conditions a) that $|hk/p|$ is sufficiently small and b) that $p/\log|k|$ is sufficiently large.
@@ 2 Notation and main results
As usual, let $\operatorname{L}^{2}(\Omega)$ denote the Lebesgue space of complex-valued square integrable functions, and define its vector-valued version ${\bf L}^{2}(\Omega):=(\operatorname{L}^{2}(\Omega))^{3}$. For a bounded Lipschitz domain $\Omega\subseteq\mathbb{R}^{3}$ and $s\geq 0$, the space $\operatorname{H}^{s}(\Omega)$ is the usual Sobolev space of order $s$, see [15, Chapter 3], and $\operatorname{H}^{s}_{0}(\Omega)$ denotes the closure of $\mathcal{C}^{\infty}_{0}(\Omega)$ in $\operatorname{H}^{s}(\Omega)$. We write $\operatorname{H}^{s}(\Omega)^{\prime}$ for the dual of $\operatorname{H}^{s}_{0}(\Omega)$ and in order to deal with vector fields, we define the vector-valued space ${\bf H}^{s}(\Omega):=(\operatorname{H}^{s}(\Omega))^{3}$. In addition, we define the spaces
@@
Finally, for a sufficiently smooth surface $\Sigma$ and $s>0$, let $\operatorname{H}^{s}(\Sigma)$ be the fractional Sobolev space of index $s$ with dual space $\operatorname{H}^{-s}(\Sigma)$, see [15, Chapter 3], and let ${\bf H}^{s}(\Sigma):=(\operatorname{H}^{s}(\Sigma))^{3}$. Furthermore, the space of square-integrable tangent fields is given by
${\bf L}^{2}_{T}(\Sigma):=\{{\bf v}\in{\bf L}^{2}(\Sigma)\ |\ {\bf v}\cdot{\bf n}=0\},$
and for $s\geq 0$ we set
@@
    The domains $\Omega,\mathcal{G}_{1},\ldots,\mathcal{G}_{n}$, are bounded and simply connected three-dimensional $\mathcal{C}^{\ell}$-domains such that $\mathcal{G}_{1},\ldots\mathcal{G}_{n}$ are mutually disjoint and satisfy $\overline{\Omega}=\overline{\mathcal{G}_{1}}\cup\ldots\cup\overline{\mathcal{G}_{n}}$.
@@
For the treatment of Maxwell’s equations with impedance boundary conditions we will need the notion of ${\boldsymbol{\mathcal{C}}}^{m}$-tensor fields acting in the tangent plane to an orientable $\mathcal{C}^{\ell}$-surface $\Sigma$. For a vector field ${\bf v}$ on $\Sigma$ we write ${\bf v}_{T}:={\bf n}\times({\bf v}\times{\bf n})$ for the tangential component of ${\bf v}$, where ${\bf n}$ is an outer unit normal to $\Sigma$.
@@ 2.2 Main results
We highlight that compared to [23, Theorem 2.2] or [5, Theorem 9.1], our result allows for inhomogeneous normal traces on $\Gamma$ and is also slightly sharper, since our estimates involve only the curl, divergence and normal trace of ${\bf u}$, but are independent of $\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)}$.
@@
$\left\|{\bf v}\right\|_{{\bf PH}^{\ell+1}(\mathcal{G})}\leq C\left\|\operatorname{curl}{\bf u}\right\|_{{\bf PH}^{\ell}(\mathcal{G})}\quad{\rm and}\quad\left\|\phi\right\|_{\operatorname{H}^{1}(\Omega)}+\left\|\phi\right\|_{\operatorname{PH}^{\ell+2}(\Omega)}\leq C\left(\left\|\operatorname{div}\boldsymbol{\nu}{\bf u}\right\|_{{\bf PH}^{\ell}(\mathcal{G})}+\left\|h\right\|_{\operatorname{H}^{\ell+1/2}(\Gamma)}\right),$
as well as $\left(\boldsymbol{\nu}{\bf v},\nabla\xi\right)_{{\bf L}^{2}(\Omega)}=0$ for all $\xi\in\operatorname{H}^{1}(\Omega)$. Moreover, the constant $C>0$ depends only on $\mathcal{G}$, $\boldsymbol{\nu}$, and $\ell$.
If, in addition, $\boldsymbol{\nu}$ satisfies $\boldsymbol{\nu}\in{\boldsymbol{\mathcal{C}}}^{\ell+1}(\overline{\Omega})$ and ${\bf u}$ is such that ${\bf u}\in{\bf H}^{\ell}(\operatorname{curl},\Omega)\cap{\bf H}^{\ell}(\operatorname{div},\Omega)$, then the above statements hold true with $\operatorname{PH}^{\ell+2}(\Omega)$, ${\bf PH}^{\ell+1}(\mathcal{G})$ and ${\bf PH}^{\ell}(\mathcal{G})$ replaced by $\operatorname{H}^{\ell+2}(\Omega)$, ${\bf H}^{\ell+1}(\Omega)$ and ${\bf H}^{\ell}(\Omega)$ in all instances.
@@
$\left\|{\bf v}\right\|_{{\bf PH}^{\ell+1}(\mathcal{G})}\leq C\left\|\operatorname{div}\boldsymbol{\nu}{\bf u}\right\|_{{\bf PH}^{\ell}(\mathcal{G})}\quad{\rm and}\quad\left\|{\bf z}\right\|_{{\bf H}^{1}(\Omega)}+\left\|\operatorname{curl}{\bf z}\right\|_{{\bf PH}^{\ell+1}(\Omega)}\leq C\left(\left\|\operatorname{curl}{\bf u}\right\|_{{\bf PH}^{\ell}(\mathcal{G})}+\left\|{\bf g}_{T}\right\|_{{\bf H}^{\ell+1/2}_{T}(\Gamma)}\right),$
as well as $\left({\bf v},\operatorname{curl}{\bf w}\right)_{{\bf L}^{2}(\Omega)}=0$ for all ${\bf w}\in{\bf H}(\operatorname{curl},\Omega)$. Moreover, the constant $C>0$ depends only on $\mathcal{G}$, $\boldsymbol{\nu}$, and $\ell$.
If, in addition, $\boldsymbol{\nu}$ satisfies $\boldsymbol{\nu}\in{\boldsymbol{\mathcal{C}}}^{\ell+1}(\overline{\Omega})$ and ${\bf u}$ is such that ${\bf u}\in{\bf H}^{\ell}(\operatorname{curl},\Omega)\cap{\bf H}^{\ell}(\operatorname{div},\Omega)$, then the above statements hold true with ${\bf PH}^{\ell+1}(\mathcal{G})$ and ${\bf PH}^{\ell}(\mathcal{G})$ replaced by ${\bf H}^{\ell+1}(\Omega)$ and ${\bf H}^{\ell}(\Omega)$ in all instances.
@@
  • ${\bf g_{\operatorname{N}}}\in{\bf H}^{\ell+1/2}_{T}(\Gamma)$ and $\operatorname{div}_{\Gamma}{\bf g_{\operatorname{N}}}\in\operatorname{H}^{\ell+1/2}(\Gamma)$, as well as ${\bf f}\cdot{\bf n}\in\operatorname{H}^{\ell+1/2}(\Gamma)$ in the case of natural boundary conditions (1.1a).
Then, there holds ${\bf u}\in{\bf PH}^{\ell+1}(\mathcal{G})\cap{\bf PH}^{\ell+1}(\operatorname{curl},\mathcal{G})$ with the estimate
$\begin{split}\left\|{\bf u}\right\|_{{\bf PH}^{\ell+1}(\mathcal{G})}+|k|^{-1}\left\|\operatorname{curl}{\bf u}\right\|_{{\bf PH}^{\ell+1}(\mathcal{G})}\leq C&\left(|k|^{\ell+1}\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)}+|k|^{\ell}\left\|\operatorname{curl}{\bf u}\right\|_{{\bf L}^{2}(\Omega)}\right)\\ &\hskip 85.35826pt+C|k|^{\ell-1}\sum_{m=0}^{\ell}\left[{\bf F}_{m,k}+{\bf G}_{m,k}\right],\end{split}$ (2.1)
where the constant $C>0$ depends only on $\mathcal{G}$, $\ell$, $\boldsymbol{\mu}^{-1}$, $\boldsymbol{\varepsilon}$, and, if necessary, $\boldsymbol{\zeta}$.
Moreover, if $\operatorname{div}{\bf f}\in\operatorname{PH}^{\ell+1}(\mathcal{G})$ and, depending on the boundary conditions, there holds that
  • $\operatorname{div}_{\Gamma}{\bf g_{\operatorname{I}}}\in\operatorname{H}^{\ell+3/2}(\Gamma)$ as well as $\boldsymbol{\zeta}$ being a coercive ${\boldsymbol{\mathcal{C}}}^{\ell+1}$-tensor field acting in the tangent plane to $\Gamma$ in the case of impedance boundary conditions (1.1a),
  • ${\bf g_{\operatorname{N}}}\in{\bf H}^{\ell+3/2}_{T}(\Gamma)$ as well as $\operatorname{div}_{\Gamma}{\bf g_{\operatorname{N}}}\in\operatorname{H}^{\ell+3/2}(\Gamma)$ and ${\bf f}\cdot{\bf n}\in\operatorname{H}^{\ell+3/2}(\Gamma)$ in the case of natural boundary conditions (1.1b),
then ${\bf u}\in{\bf PH}^{\ell+2}(\mathcal{G})$, and in addition to (2.1) there holds
$\begin{split}\left\|{\bf u}\right\|_{{\bf PH}^{\ell+2}(\mathcal{G})}\leq C&\left(|k|^{\ell+1}\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)}+|k|^{\ell}\left\|\operatorname{curl}{\bf u}\right\|_{{\bf L}^{2}(\Omega)}+|k|^{\ell-1}\left\|\operatorname{div}{\bf f}\right\|_{\operatorname{PH}^{\ell+1}(\mathcal{G})}+|k|^{\ell-1}{\bf G}_{\ell+1,k}\right)\\ &\hskip 213.39566pt+C|k|^{\ell-1}\sum_{m=0}^{\ell}\left[{\bf F}_{m,k}+{\bf G}_{m,k}\right],\end{split}$ (2.2)
where the constant $C>0$ again depends only on $\mathcal{G}$, $\ell$, $\boldsymbol{\mu}^{-1}$, $\boldsymbol{\varepsilon}$, and, if necessary, $\boldsymbol{\zeta}$.
Let us remark that in general there are $k\in\mathbb{C}\setminus\{0\}$, right-hand sides ${\bf f}$ and, if necessary, boundary data ${\bf g_{\operatorname{I}}}$ or ${\bf g_{\operatorname{N}}}$ such that no weak solution ${\bf u}$ of (1.1) exists. However, if there is a (possible not unique) weak solution ${\bf u}$, then Theorem 2.7 asserts piecewise Sobolev regularity of any weak solution of (1.1).
@@
  • in the case of impedance boundary conditions (1.1a), let $\boldsymbol{\zeta}$ be a coercive ${\boldsymbol{\mathcal{C}}}^{\infty}$-tensor field acting in the tangent plane to $\Gamma$ and let ${\bf g_{\operatorname{I}}}$ be a smooth tangent field on $\Gamma$,

### v3

@@ Abstract
We consider a bounded Lipschitz domain $\Omega\subseteq\mathbb{R}^{3}$ with sufficiently smooth boundary and prove piecewise Sobolev regularity of vector fields that have piecewise regular curl and divergence, but may be discontinuous across mutually disjoint and sufficiently smooth surfaces inside of $\Omega$. The main idea behind our approach is to employ recently developed parametrices for the curl-operator and the regularity theory of Poisson transmission problems. We conclude our work by applying our findings to the heterogeneous time-harmonic Maxwell equations with either a) impedance, b) natural or c) essential boundary conditions and providing wavenumber-explicit piecewise regularity estimates for these equations.
@@ 1 Introduction
[…]
@@
Let us mention that this work is the first paper of an upcoming series of research articles. The second paper [18] in this series builds on the subsequent Theorem 2.6 and Theorem 2.7 and proves wavenumber-explicit piecewise analyticity of solutions $\bf u$} of (1.1), provided that the geometry and all given coefficients and data are (piecewise) analytic. Subsequently, the third work [19] of this upcoming series provides a wavenumber-explicit analysis of the $hp$-finite element method applied to Maxwell’s equations (1.1) in the presence of piecewise analytic coefficients. In particular, if $k$ is the wavenumber, $h$ is the local mesh-width and $p$ is the local polynomial degree in the $hp$-FEM approximation, the third paper will show that the $hp$-FEM is quasi-optimal under the scale-resolution conditions a) that $|hk/p|$ is sufficiently small and b) that $p/\log|k|$ is sufficiently large.
@@ 2 Notation and main results
As usual, let $\operatorname{L}^{2}(\Omega)$ denote the Lebesgue space of complex-valued square integrable functions, and define its vector-valued version ${\bf L}^{2}(\Omega):=(\operatorname{L}^{2}(\Omega))^{3}$. For a bounded Lipschitz domain $\Omega\subseteq\mathbb{R}^{3}$ and $s\geq 0$, the space $\operatorname{H}^{s}(\Omega)$ is the usual Sobolev space of order $s$, see [15, Chapter 3], and $\operatorname{H}^{s}_{0}(\Omega)$ denotes the closure of $\mathcal{C}^{\infty}_{0}(\Omega)$ in $\operatorname{H}^{s}(\Omega)$. Furthermore, in order to deal with vector fields, we define the vector-valued space ${\bf H}^{s}(\Omega):=(\operatorname{H}^{s}(\Omega))^{3}$ and set
@@
Finally, for a sufficiently smooth and orientable surface $\Sigma$ and $s>0$, let $\operatorname{H}^{s}(\Sigma)$ be the fractional Sobolev space of index $s$ with dual space $\operatorname{H}^{-s}(\Sigma)$, see [15, Chapter 3], and let ${\bf H}^{s}(\Sigma):=(\operatorname{H}^{s}(\Sigma))^{3}$. Furthermore, since $\Sigma$ is supposed to be orientable we may choose a normal unit vector field $\bf n$} on $\Sigma$ and define the space of square-integrable tangent fields by
${\bf L}^{2}_{T}(\Sigma):=\{{\bf v}\in{\bf L}^{2}(\Sigma)\ |\ {\bf v}\cdot{\bf n}=0\}.$
For $s\geq 0$ we set
@@
    The domains $\Omega,\mathcal{G}_{1},\ldots,\mathcal{G}_{n}$ are bounded three-dimensional $\mathcal{C}^{\ell}$-domains and $\Omega$ is simply connected. Furthermore, the domains $\mathcal{G}_{1},\ldots\mathcal{G}_{n}$ are mutually disjoint and satisfy $\overline{\Omega}=\overline{\mathcal{G}_{1}}\cup\ldots\cup\overline{\mathcal{G}_{n}}$.
@@
Remark 2.2.
The assumption that $\Omega$ is simply connected makes some proofs in this work simpler (especially the proofs of Lemma 4.5 and Lemma 4.6), but it is not fundamental for our results. Indeed, the regularity results of Theorem 2.6 and Theorem 2.7 below hold true for a non-simply connected domain $\Omega$ as well with only minor modifications necessary. The difference between the simply connected case considered in the following and the non-simply connected case is that in the latter the cohomology spaces of the de Rham complex have to be taken into account. It is known that on a bounded Lipschitz domain, these cohomology spaces are finite dimensional spaces of smooth functions or vector fields [10]. Hence, the existence of non-trivial cohomology spaces does not spoil the regularity results that we derive for the case of simply connected $\Omega,\mathcal{G}_{1},\ldots,\mathcal{G}_{n}$, only the norm bounds have to be adapted. We come back to this issue in Remark 2.9 and Remark 2.11.
@@
For the treatment of Maxwell’s equations with impedance boundary conditions we will need the notion of ${\boldsymbol{\mathcal{C}}}^{m}$-tensor fields acting in the tangent plane to an orientable $\mathcal{C}^{\ell}$-surface $\Sigma$. For a vector field $\bf v$} on $\Sigma$ we write ${\bf v}_{T}:={\bf n}\times({\bf v}\times{\bf n})$ for the tangential component of $\bf v$}, where $\bf n$} is a unit normal to $\Sigma$.
@@
Notice that we defined ${\bf PH}^{m}(\operatorname{curl},\mathcal{G})$ and ${\bf PH}^{m}(\operatorname{div},\mathcal{G})$ in a rather unconventional way. Canonically, one would rather write ${\bf v}\in{\bf PH}^{m}(\operatorname{curl},\mathcal{G})$ if ${\bf v}\in{\bf PH}^{m}(\mathcal{G})$ and $\operatorname{curl}{\bf v}\in{\bf PH}^{m}(\mathcal{G})$ and analogously for ${\bf PH}^{m}(\operatorname{div},\mathcal{G})$. We deliberately chose this unorthodox way in order to be able to talk about (piecewise) regularity of $\operatorname{curl}{\bf u}$ and $\operatorname{div}{\bf u}$ independently of the (piecewise) regularity of $\bf u$}.
@@ 2.2 Main results
We highlight that compared to [23, Theorem 2.2] or [5, Theorem 9.1], our result allows for inhomogeneous normal traces on $\Gamma$ and is also slightly sharper, since our estimates involve only the curl, divergence and normal trace of $\bf u$}, but are independent of $\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)}$. This independence of $\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)}$ is due to our assumptions of $\Omega$ being simply connected and of $\Gamma$ and $\mathcal{I}$ consisting of simply connected components.
@@
$\begin{split}\left\|{\bf v}\right\|_{{\bf PH}^{\ell+1}(\mathcal{G})}&\leq C\left\|\operatorname{curl}{\bf u}\right\|_{{\bf PH}^{\ell}(\mathcal{G})},\\ \left\|\phi\right\|_{\operatorname{H}^{1}(\Omega)}+\left\|\phi\right\|_{\operatorname{PH}^{\ell+2}(\Omega)}&\leq C\left(\left\|\operatorname{div}\boldsymbol{\nu}{\bf u}\right\|_{{\bf PH}^{\ell}(\mathcal{G})}+\left\|h\right\|_{\operatorname{H}^{\ell+1/2}(\Gamma)}\right),\end{split}$ (2.1)
and $\left(\boldsymbol{\nu}{\bf v},\nabla\xi\right)_{{\bf L}^{2}(\Omega)}=0$ for all $\xi\in\operatorname{H}^{1}(\Omega)$. Moreover, the constant $C>0$ depends only on $\mathcal{G}$, $\boldsymbol{\nu}$, and $\ell$.
@@
$\begin{split}\left\|{\bf v}\right\|_{{\bf PH}^{\ell+1}(\mathcal{G})}&\leq C\left\|\operatorname{div}\boldsymbol{\nu}{\bf u}\right\|_{{\bf PH}^{\ell}(\mathcal{G})},\\ \left\|{\bf z}\right\|_{{\bf H}^{1}(\Omega)}+\left\|\operatorname{curl}{\bf z}\right\|_{{\bf PH}^{\ell+1}(\Omega)}&\leq C\left(\left\|\operatorname{curl}{\bf u}\right\|_{{\bf PH}^{\ell}(\mathcal{G})}+\left\|{\bf g}_{T}\right\|_{{\bf H}^{\ell+1/2}_{T}(\Gamma)}\right),\end{split}$ (2.2)
and $\left({\bf v},\operatorname{curl}{\bf w}\right)_{{\bf L}^{2}(\Omega)}=0$ for all ${\bf w}\in{\bf H}(\operatorname{curl},\Omega)$. Moreover, the constant $C>0$ depends only on $\mathcal{G}$, $\boldsymbol{\nu}$, and $\ell$.
Remark 2.8.
In the case that the $\mathcal{C}^{\ell+2}$-partition $\mathcal{G}$ consists only of one subdomain (that is, there are no surfaces of discontinuity of $\boldsymbol{\nu}$), the statements of Theorem 2.6 and Theorem 2.7 hold with the broken Sobolev norms replaced by their ”usual” counterparts: If $\boldsymbol{\nu}$ satisfies $\boldsymbol{\nu}\in{\boldsymbol{\mathcal{C}}}^{\ell+1}(\overline{\Omega})$ and ${\bf u}\in{\bf H}^{\ell}(\operatorname{curl},\Omega)\cap{\bf H}^{\ell}(\operatorname{div},\Omega)$, then the statements of Theorem 2.6 and Theorem 2.7 hold true with ${\bf PH}^{\ell+1}(\mathcal{G})$ and ${\bf PH}^{\ell}(\mathcal{G})$ replaced by ${\bf H}^{\ell+1}(\Omega)$ and ${\bf H}^{\ell}(\Omega)$ in all instances.
Remark 2.9.
In the case that $\Omega$ is not simply connected or if $\Gamma$ or the subdomain interface $\mathcal{I}$ have non-simply connected components, Theorem 2.6 and Theorem 2.7 still remain valid once we change (2.1) and (2.2) to
$\left\|{\bf v}\right\|_{{\bf PH}^{\ell+1}(\mathcal{G})}$ $\leq C\left(\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)}+\left\|\operatorname{curl}{\bf u}\right\|_{{\bf PH}^{\ell}(\mathcal{G})}\right),$
$\left\|\phi\right\|_{\operatorname{H}^{1}(\Omega)}+\left\|\phi\right\|_{\operatorname{PH}^{\ell+2}(\Omega)}$ $\leq C\left(\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)}+\left\|\operatorname{div}\boldsymbol{\nu}{\bf u}\right\|_{{\bf PH}^{\ell}(\mathcal{G})}+\left\|h\right\|_{\operatorname{H}^{\ell+1/2}(\Gamma)}\right),$
and
$\left\|{\bf v}\right\|_{{\bf PH}^{\ell+1}(\mathcal{G})}$ $\leq C\left(\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)}+\left\|\operatorname{div}\boldsymbol{\nu}{\bf u}\right\|_{{\bf PH}^{\ell}(\mathcal{G})}\right),$
$\left\|{\bf z}\right\|_{{\bf H}^{1}(\Omega)}+\left\|\operatorname{curl}{\bf z}\right\|_{{\bf PH}^{\ell+1}(\Omega)}$ $\leq C\left(\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)}+\left\|\operatorname{curl}{\bf u}\right\|_{{\bf PH}^{\ell}(\mathcal{G})}+\left\|{\bf g}_{T}\right\|_{{\bf H}^{\ell+1/2}_{T}(\Gamma)}\right),$
respectively, cf. Remark 2.2.
@@
  • ${\bf g_{\operatorname{N}}}\in{\bf H}^{\ell+1/2}_{T}(\Gamma)$ and ${{\bf f}\cdot{\bf n}-}\operatorname{div}_{\Gamma}{\bf g_{\operatorname{N}}}\in\operatorname{H}^{\ell+1/2}(\Gamma)$ in the case of natural boundary conditions (1.1b).
Then, for all $m\in\{0,\ldots,\ell\}$ there holds ${\bf u}\in{\bf PH}^{{m}+1}(\mathcal{G})\cap{\bf PH}^{{m}+1}(\operatorname{curl},\mathcal{G})$ with the estimate
$\begin{split}{|k|^{-m-1}}\left\|{\bf u}\right\|_{{\bf PH}^{{m}+1}(\mathcal{G})}+{|k|^{-m-2}}\left\|\operatorname{curl}{\bf u}\right\|_{{\bf PH}^{{m}+1}(\mathcal{G})}\leq{C_{m}}&\left(\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)}+|k|^{{-1}}\left\|\operatorname{curl}{\bf u}\right\|_{{\bf L}^{2}(\Omega)}\right)\\ &\hskip 85.35826pt+{C_{m}}|k|^{{-2}}\sum_{j=0}^{{m}}\left[{\bf F}_{j,k}+{\bf G}_{j,k}\right],\end{split}$ (2.3)
where the constant ${C_{m}}>0$ depends only on $\mathcal{G}$, m{m}, $\boldsymbol{\mu}^{-1}$, $\boldsymbol{\varepsilon}$, and, if necessary, $\boldsymbol{\zeta}$.
Remark 2.11.
If $\Omega$ is not simply connected or if $\Gamma$ or the subdomain interface $\mathcal{I}$ have non-simply connected components, Theorem 2.10 still remains valid once we change (2.3) to
$\begin{split}{|k|^{-m-1}}\left\|{\bf u}\right\|_{{\bf PH}^{{m}+1}(\mathcal{G})}+{|k|^{-m-2}}\left\|\operatorname{curl}{\bf u}\right\|_{{\bf PH}^{{m}+1}(\mathcal{G})}\leq{C_{m}}&\left(1+\sum_{j=0}^{m}|k|^{-j-1}\right)\left(\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)}+|k|^{{-1}}\left\|\operatorname{curl}{\bf u}\right\|_{{\bf L}^{2}(\Omega)}\right)\\ &\hskip 85.35826pt+C_{m}|k|^{{-2}}\sum_{j=0}^{{m}}\left[{\bf F}_{j,k}+{\bf G}_{j,k}\right].\end{split}$
For $|k|\geq k_{0}>0$ this reduces again to (2.3) for a modified constant $C_{m}$ that depends also on $k_{0}$.
Let us highlight that in general there are $k\in\mathbb{C}\setminus\{0\}$, right-hand sides $\bf f$} and, if necessary, boundary data $\bf g_{\operatorname{I}}$} or $\bf g_{\operatorname{N}}$} such that no weak solution $\bf u$} of (1.1) exists. However, if there is a (possible not unique) weak solution $\bf u$}, then Theorem 2.10 asserts piecewise Sobolev regularity of any weak solution of (1.1).
@@
[…]